OpenStudy (anonymous):

so the question was equation of a slope 4/5 passing thru the points (-1,3)...the answer I got was y=4/5x +11. how far off am I??

OpenStudy (phi):

point slope form is \[ y - y_0 = m(x-x_0 )\] where \( (x_0,y_0) \) is a point on the line, and m is the slope here: \[ y - 3 = \frac{4}{5}(x- -1) \\ y -3 = \frac{4}{5}(x+1)\] if you want this in slope-intercept form, distribute the 4/5 (multiply each term inside the parens by 4/5) : \[ y -3 = \frac{4}{5}x+\frac{4}{5}\] now add +3 to both sides and simplify

OpenStudy (anonymous):

very helpful! thanks!

OpenStudy (phi):

btw, we will not get y=4/5x +11

OpenStudy (anonymous):

yes...i saw why! thanks again!!